Simplify the equation y= 1/4log(base of 2)(8x-56)^16 -12 by applying law of logarithms? How are the original equation and the simplified equation related?
Answer by stanbon(75887)

Simplify the equation y= 1/4log(base of 2)(8x-56)^16 -12 by applying law of logarithms? How are the original equation and the simplified equation related?
------
y = (1/4)log2[8(x-7)] - 12
---
y = (1/4)[log2(8)+log2(x-7)] - 12
---
y = (1/4)[3+log2(x-7)] - 12
--------------------
Cheers,
Stan H.
-----------